Please help me answer this question completely ( please don't answer if not complete), better by typing, if handwritten please write clearly. Thank you.

Subject: Inorganic Chemistry

Question: Below are polynuclear complexes. CALCULATE the numbers of metal-metal (M-M) bond and then , ELUCIDATE the structure.

(a) Cp2 Ni2Fe (M3-S) (CO)3 (b) [CpMo(CO)2]2
Answer #1

And every metal atom have satisfied the 18e - rule Ni Cp 10+ 5 Ni + 1+ s Fe 1+1 1 - =18 Ž 2 Fe same Fe 300s Ni, No 1310 The o6) Similarly, for the compound [ep Mo (Co),] total no. of valence electrons are 30 (cp=5+ M0 = 6 + 2.00 =4)] Hence total no.a) CP₂ Niz Fe (M3-3) (CO₂ b) [Cp Mo (Co), I 7 a) C. p. Nig Fe (M3-5) (CO). According to 18 e rule the number of electron counThe total numbers of metal-metal bond for both of the complex is 3 and the calculation and the structure is given in the attached images.